can the system
3xy-4y-z=5
3x+2y-z=0
x-y+z=1
be solved by using Cramer's rule? if not, why? if yes, solve it.
1
Expert's answer
2020-05-06T18:24:22-0400

"\begin{vmatrix}
 3 & -4 & -1 \\
 3 & 2 & -1 \\
 1 & -1 &1
\end{vmatrix} =6+4+3+2+12-3=24"


"\begin{vmatrix}
 5 & -4 & -1 \\
 0 & 2 & -1 \\
 1 & -1 &1
\end{vmatrix} =10+4+0+2+0-5=11"


"\begin{vmatrix}
 3 & 5 & -1 \\
 3 & 0 & -1 \\
 1 & 1 &1
\end{vmatrix} =0-3-5+0-15+3=-20"


"\begin{vmatrix}
 3 & -4 & 5 \\
 3 & 2 & 0 \\
 1 & -1 &1
\end{vmatrix} =6-15+0-10+12-0=-7"


Hence, "x=\frac{11}{24}, y=-\frac{20}{24}, z=-\frac{7}{24}"